Status Report

ISS On-Orbit Status 13 Jan 2003

By SpaceRef Editor
January 13, 2003
Filed under , ,
ISS On-Orbit Status 13 Jan 2003

All ISS systems continue to function nominally, except as noted previously or below.  Begin of Week 7 for Expedition 6, with an exciting spacewalk event on Wednesday (1/15).

In preparation for the EVA, CDR Ken Bowersox and FE-2/SO Don Pettit completed another session with the medical assessment MO-9 of the Russian crew health-monitoring program right after wakeup (1:00am EST) prior to breakfast and before physical exercise.  Later, they finished the clinical U.S. PHS (periodic health status) exam, without blood labs.  [MO-9 is biochemical urinalysis, conducted regularly every 30 days (and sometimes also before and after EVAs) with the Urolux equipment.  It is one of five nominal Russian medical tests that were adopted by NASA for US crewmembers for the PHS evaluation.  PHS, guided by software on the MEC (medical equipment computer), is performed once every 30 days for each crewmember and two weeks before landing and as clinically indicated.  PHS evaluations alternate between the one with blood labs and the one without.  The tests are documented afterwards, and physical exercise should then follow.]

Meanwhile, FE-1 Nikolai Budarin conducted his regular daily inspection of the Russian BIO-5 Rasteniya-2/Lada-2 (“Plants-2”) plant growth experiment.

Budarin verified two communications configurations for the EVA in the Service Module (SM).  [This task consisted, first, of checking out comm setup 2B, deactivating SM comm panel 3 to ensure no echo is heard at MCC-Moscow and MCC-Houston during the spacewalk, and verifying that VHF-2 was off and channel 3 was selected on comm panel 2.  Then, for comm configuration 1A, he selected channel 1 on comm panel 3, verified VHF 2 was off and selected channel 3 on panel 2.]

Configuration of the communications and tracking systems for and after the EVA was commanded by the ground.
After central ventilation fans TsV-1 and -2 in the FGB were deactivated by ground control, Bowersox spent about an hour cleaning the protective mesh screens (grills) of the ventilators.  Later, the fans were commanded back on.

Bowersox and Pettit completed the planned EVA dry-run activities, except for SAFER (simplified aid for EVA rescue) checkout.  [First, Airlock (A/L) depress operations to 10.2 psi and Node LTL RFCA (low temperature loop/rack flow control assembly) configuration were initiated by the ground, to support the EMU (extravehicular mobility unit) donning dry-run (with suit pressurization and fit verification, EMU TV test, heated glove checkout).  Next, EMU prebreathe protocol was  exercised, followed by post-EVA activities (suit doffing and powerdown).  The crew used the same procedures that will be used for the spacewalk on Wednesday, but many steps were only simulated for this event.  No O2 (oxygen) flowed through the PHAs (prebreathe hose assemblies) during the prebreathe exercise on the CEVIS cycle ergometer.  The PHA masks, per ground instruction, were replaced by the quick-don PBA masks (portable breathing apparatus).  Water recharge for the suits was not required afterwards, nor was Metox regeneration.  10.2-psi operations in the A/L were then terminated.]

CDR Bowersox completed the daily routine tasks of SOSh life support systems maintenance and Lab payload status checkup (PCG-STES010, ZCG), while FE-2 Pettit prepared the daily IMS inventory delta file for automated export to MCC-M.

FE-1 Budarin conducted an inspection of the BRPK air/condensate separator system which he had worked on yesterday.  His replacement of the SRVK (lane 1)/BRPK-1 appears to have been successful.

Later, Nikolai completed the periodic visual inspection and photo documentation of the SM and DC-1 (docking compartment) windows.  [Objective of the inspection, using digital still camera and voice recorder, is to assess the external panes of the windows for any changes (new cavities, scratches, new or expanded old stains or discolorations effecting transparency properties).  The last inspection was performed by the Expedition 5 crew, and the new assessment is to be compared to the earlier observations.  Photo file preparation and downlink of the images is scheduled for tomorrow.]

The planned CCAA (common cabin air assembly) swap from starboard to port air conditioner was successfully performed yesterday.  The temperature setpoint was set to 21.2 degC.  The crew is free to continue to set the temperature, giving ECLSS (environment control and life support system) specialists the usual heads-up.

The Vozdukh CO2 scrubber is reportedly failed.  Troubleshooting is expected tomorrow.  The U.S. CDRA  (CO2 removal assembly) is operating.

Tasks completed by the crew over last weekend included A/L E/L (equipment lock) preparation, FFQ (food frequency questionnaire) logging for Bowersox, and the periodic PCS (portable computer system) reboots.

Today’s targets for the CEO (crew earth observations program) were South Island, New Zealand (Dynamic Event:  Clear weather over South Island after many weeks of no/low visibility.  Dramatic pass over the Southern Alps with low sun angles), Bamako, Mali (nadir pass; ESC [electronic still camera]), West African fires (Dynamic Event:  fires rage ever more numerous and widespread in the tree-grass savannas of West Africa), Angolan Biomass Burning (although this is the low-fire [wet] season, some fires may exist, and burn scars from previous seasons are of interest [as irregular shapes often superimposed, so that the most recent fires are the clearest, several fire seasons being recorded in fire-prone areas]), and Industrialized SE Africa (high pressure should have allowed visible accumulation of aerosols [industrial mainly, biomass burning and dust related].  Crew was to shoot obliques right.  Also, near nadir views of southern Lesotho mountains just left).
CEO images can be viewed at the website

U.S. and Russian Segment Status (as of 1:35pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On (16 Amp mode).  Vozdukh CO2 scrubber is Off.  U.S. CDRA CO2 scrubber is On. TCCS (trace contaminant control subsystem) is operational.  MCA (major constituents analyzer) is failed.  BMP Harmful Impurities unit: Absorbent bed #1 in Purify mode, bed #2 in Purify mode.
  • SM Working Compartment:  Pressure (mmHg) — 736; temperature (deg C) — 22.5; ppO2 (mmHg) — 154.7; ppCO2 (mmHg) — 3.8 (suspect).
  • SM Transfer Compartment:  Pressure (mmHg) — 741; temperature (deg C) — 18.1.
  • FGB Cabin:  Pressure (mmHg) — 744; temperature (deg C) — 20.3.
  • Node:  Pressure (mmHg) — 734.98; temperature (deg C) — 22.8 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 737.34; temperature (deg C) — 23.2; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;
  • Joint Airlock (Equip. Lock):  Pressure (mmHg) — 737.44, temperature (deg C) — n/a; shell heater temp (deg C) — 22.0, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1:  Shell heater temp (deg C) — 22.9
  • PMA-2:  Shell heater temp (deg C) — 19.3

(n/a = data not available)

Propulsion System (PS): 

  • Total propellant load available (SM + FGB + Progress) — 3689 kg (8133 lb) [as of 1/9/03].

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  Beta Gimbal Assembly (BGA) 2B and BGA 4B both in Dual-angle mode (directed position).
  • SM batteries (3am):  Battery #7 is off line (failed); battery #3 is in “Cycle” mode; all other batteries (6) are in “Partial Charge” mode.
  • FGB batteries (3am):  Batteries #3 is offline (failed); battery #6 is in “Cycle” mode; all other batteries (4) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 is in Standby; PCU-2 is in Standby.

Thermal Control Systems:

  • Air conditioner SKV-1 is On, SKV-2 is Off.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup (new patches loaded on both).
  • EXT-2 is On (primary), EXT-1 is off.
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is On (primary); PL-2 MDM is off (diagnostic
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational. Lane 1 is down (as of 11/14).
  • SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.

Attitude Source:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector — U.S. SIGI-1
  • Attitude — U.S. SIGI-1
  • Angular rates — U.S. RGA-1

Flight Attitude:

  • LVLH TEA (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch: -9.1 deg, roll: 0 deg]).
  • Solar Beta Angle: 12.7 (magnitude increasing)

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operating.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally.
  • Ku-band is operating nominally.
  • Audio subsystem operating nominally.
  • Video subsystem operating nominally.
  • MCOR (medium-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings.  MT: at WS7, with KA power.  POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is Off; Cupola RWS is Off.

ISS Orbit (as of this morning, 7:26am EST [= epoch]):

  • Mean altitude — 390.5 km
  • Apogee — 394.0 km
  • Perigee — 387.0 km
  • Period — 92.37 min.
  • Inclination (to Equator) — 51.63 deg
  • Eccentricity — 0.0005175
  • Orbits per 24-hr. day — 15.59
  • Solar Beta Angle — 12.7 deg (magnitude increasing)
  • Altitude loss — 210 m (mean) in last 24 hours
  • Revolutions since FGB/Zarya launch (Nov. ’98) — 23687

For more on ISS orbit and worldwide naked-eye visibility dates/times, see

SpaceRef staff editor.